OpenStudy (anonymous):

An airplane can fly at a rate of 560 mph in calm air. It can fly 2400 miles with the wind at the same time it can fly 2080 miles against the wind. Solve for the speed of the wind.

OpenStudy (dumbcow):

d = r*t 2400 = (560+x)*t 2080 = (560-x)*t solve for t, set equations equal --> 2400/(560+x) = 2080/(560-x) --> 2400(560-x) = 2080(560+x) --> 4480x = 179,200 --> x = 40